# SPLITZ: Certifiable Robustness via Split Lipschitz Randomized Smoothing thanks: This work was supported by NSF grants CCF 2100013, CNS 2209951, CCF 1651492, CNS 2317192 and CNS 1822071. This wrok has been submitted to IEEE Transactions on Information Forensics and Security

### v1

@@ IV Experiments and Evaluation Results
In this section, we evaluate the SPLITZ classifier on two datasets, MNIST[36] and CIFAR-10 [37], where we demonstrate that our proposed approach consistently surpasses other state-of-the-art methods. For all datasets, we report the approximate certified test accuracy and certified radius of smoothed classifiers over test datasets (full test datasets in MNIST and CIFAR-10 datasets). Same as previous works, we vary the noise level $\sigma\in\{0.25,0.5,1.0\}$ for the smoothed models on CIFAR-10 dataset ($\sigma\in\{0.25,0.5,0.75,1.0\}$ for MNIST dataset) and certified the same noise level $\sigma$ during the inference time. To ensure a fair comparison with previous studies, we provide the highest reported results from each paper for the corresponding above levels of noise magnitudes. To improve certified robustness, we utilize the tighter local Lipschitz bound introduced in [19]. For two datasets, we use the same model as previous works [7, 14, 12, 22] (LeNet for MNIST, ResNet110 for CIFAR-10). More experimental details are described in Appendix C.
@@
TABLE II: Comparison of the approximate certified test accuracy (%) on CIFAR-10 under $\ell_{2}$ norm perturbation. Extra data indicates whether their models incorporate other datasets in their models. Each entry lists the certified accuracy using numbers taken from respective papers. We set our values bold face when the value outperforms the best among the comparison methods.
Certified accuracy at $\epsilon~{}(\%)$
@@
SPLITZ Methodology For all datasets, we split the classifier after $1^{\text{st}}$ affine layer where the left half of the classifier contains one convolution layer followed by the clipped ReLU layer (See Appendix B). The rest of the classifier is the same as original models (LeNet for MNIST, ResNet110 for CIFAR-10).
Dataset Configuration For both the MNIST and CIFAR-10 dataset(s), we draw $n_{0}=10^{2},~{}n_{1}=10^{5}$ noise samples to certify the smoothing model following [7, 14, 12]. We set the Lipschitz threshold (see Sec III) as $\theta=0.5$. For local Lipschitz constrained training, we set tradeoff parameter $\lambda$ (see Sec III) evenly decrease from $0.8-0.4$, $0.7-0.5$ respectively. We use one Nvidia P100 GPU to train the SPLITZ model with batch size $512,256$ respectively. We apply Adam Optimizer for two datasets (MNIST, CIFAR-10). For the MNIST dataset, we conduct training for 150 epochs and utilize an early-stop strategy to search for the optimal classifier over an additional 150 epochs. We set the initial learning rate as 0.001. The learning rate is decayed (multiplied by 0.1) by 0.1 at every 50 epochs (50th, 100th…). For the CIFAR-10 dataset, we train 200 epochs using the ResNet110 and use the early-stop strategy to search for the optimal classifier over an additional 200 epochs. Furthermore, we set the initial learning rate as 0.001 and final learning rate as $10^{-6}$. The learning rate starts to evenly decay at each epoch from the half of the training epochs. We report our more experimental details in Appendix C.
@@ IV-A Main Results
Results on MNIST As showed in Table I, we can observe that SPLITZ outperforms other state-of-art approaches in almost every value of $\epsilon$. Impressively, we find that the SPLITZ classifier has a significant improvement when the value of $\epsilon$ is large. For instance, when $\epsilon=2.50$, SPLITZ classifier achieves 40.5% compared to state-of-art top-1 test accuracy 34.8% certified test accuracy on the MNIST dataset. Moreover, when we increase $\epsilon$ from 1.50 to 2.50, RS drops from 67.3% to 10.9% decreasing 56.4% test accuracy. SPLITZ, however, maintains higher certified test accuracy from 80.2% to 40.5% maintaining relatively higher test accuracy.
@@
Results on CIFAR10 As shown in Table II and Fig 3, our method outperforms the state-of-art approaches for every value of $\epsilon$ on CIFAR-10 dataset. Interestingly, we find that the SPLITZ training has a significant improvement when the value of $\epsilon$ is large. For instance, when $\epsilon=1.0$, the model achieves 43.2% top-1 test accuracy on CIFAR-10 dataset compared to state-of-art top-1 test accuracy 39.8%. One hypothesis is that minimizing the Lipscitz bound of $f_{L}$ ($L_{f_{L}}\leq 1$) is able to boost the certified radius of the model. Intuitively, more samples are correctly classified while corresponding radius are larger than given $\epsilon$. In addition, we can observe the similar trend as MNIST dataset. SPLITZ maintains higher certified test accuracy when we increase $\epsilon$ from 0.25 to 1.00 compared to other state-of-art mechanisms.
@@
TABLE III: Comparison of average certified radius (ACR) across two different datasets (MNIST, CIFAR-10). We can observe that for two datasets, SPLITZ consistently achieves better results compared to other state-of-art mechanisms. * is reported by [22, 12]
$\sigma$ Methods MNIST CIFAR-10
0.25 $\text{RS}^{*}$ 0.911 0.424
$\text{SmoothAdv}^{*}$ 0.932 0.544
MACER 0.918 0.556
Consistency 0.928 0.552
SmoothMix 0.933 0.548
SPLITZ 1.752 0.680
0.50 $\text{RS}^{*}$ 1.553 0.525
$\text{SmoothAdv}^{*}$ 1.687 0.684
MACER 1.583 0.726
Consistency 1.697 0.726
SmoothMix 1.694 0.737
SPLITZ 2.059 0.924
1.00 $\text{RS}^{*}$ 1.620 0.542
SmoothAdv 1.779 0.660
MACER 1.520 0.792
Consistency 1.819 0.816
SmoothMix 1.823 0.773
SPLITZ 2.104 0.979
@@
Results on Average Certified Radius (ACR) We investigate the performance of SPLITZ using average certified radius (ACR), where we measure the correct samples’ average certified radius over the test datasets (MNIST, CIFAR-10). As shown in Table III, we provide the comprehensive comparison results of average certified radius (ACR) compared to other certified robust techniques. Our SPLITZ consistently outperforms others. For instance, when $\sigma=0.25~{}(0.50)$, the ACR of SPLITZ is 1.752 (2.059) respectively, where the state-of-the-art is 1.697 (0.933) respectively. In addition, we can observe the same trends in CIFAR-10 dataset, where SPLITZ consistently outperforms the state-of-the-art when $\sigma=0.25,0.50$ and $1.00$.
TABLE IV: SPLITZ training time and certifying time across two datasets (MNIST using LeNet, CIFAR-10 using ResNet 110 ).
Training
Datasets Architecture # of GPUs Training per epoch (s) Certifying per image (s)
MNIST LeNet 1 15.1 1.0
CIFAR-10 ResNet110 1 59.5 50.7
TABLE V: Comparison of certified test accuracy of SPLITZ with Gaussian noise $\sigma=0.75$ for varying the splitting layer on MNIST dataset with LeNet.
Certified Test Accuracy at $\epsilon~{}(\%)$
Location of Splitting 0.50 0.75 1.00 1.25 1.50 1.75 2.00 2.25 2.50
$1^{\text{st}}$ affine layer 94.1 92.0 88.8 84.7 79.0 71.3 62.3 51.7 40.5
$2^{\text{nd}}$ affine layer 89.7 86.2 81.0 75.0 68.0 59.9 51.0 41.4 31.8
$3^{\text{rd}}$ affine layer 84.4 80.4 75.6 69.9 63.6 56.9 49.4 40.3 29.3
@@ IV-B Training and certifying time
Our SPLITZ model needs to vary the value of $\gamma$ (the size of the ball around input $x$) during the training epoch. Thus, we may need relatively more time to obtain the optimal model. To solve this, we apply the early stop mechanism to obtain the better optimized model during the training process. At the same time, we decay our learning rates during the training process. We use one Nvidia P100 GPU to train our SPLITZ model and report our training time (certifying time) for each one epoch (one image) in the Table IV.
@@ IV-C Ablation Study
Impact of splitting location As mentioned in Section III, our SPLITZ classifier can be optimized over different split ways, where we conduct the experiments and show our results in Table V. For example, when $\epsilon=2$, splitting after the $1^{st}$, or $2^{nd}$, or $3^{rd}$ layer result in certified accuracy of $62.3$%, $51.0$% and $49.4$% respectively. These results indicate that splitting the neural network early achieves better performance. Intuitively, splitting the neural network early helps the model minimize the local Lipschitz bound, which improves the certified robustness leading to a higher certified test accuracy given the same $\epsilon$. As the splitting becomes ”deeper”, estimating the local Lipshitz constant also becomes harder, which implies that a looser bound leads to smaller certified radius.
@@
Comparison of the orthogonal neural network (OrthoNN) and the SPLITZ In the baseline RS+OrthoNN model, during the training process, we enforce the first convolution layer of the classifier’s left half to act as an orthogonal convolution neural network. This is achieved by incorporating an orthogonal constraint regularization into the loss function. For the SPLITZ model, we impose a constraint on the local Lipschitz constant of the classifier’s left half. The comparison between RS+orthoNN and SPLITZ is illustrated in Table I with respect to the MNIST dataset and in Table II with respect to the CIFAR-10 dataset. Our observations indicate that enforcing an orthogonal convolution layer on the classifier’s left half has a comparatively lesser impact than training with a local Lipschitz constraint.
TABLE VI: Comparison of certified test accuracy of SPLITZ with Gaussian noise $\sigma=0.75$ for varying $Lip_{\theta}$ (the threshold of the local Lipschitz bound around input $x$) on MNIST dataset.
@@
TABLE VII: Comparison of certified test accuracy of SPLITZ with Gaussian noise $\sigma=0.75$ for varying $\gamma$ (the size of the ball around input $x$) on MNIST dataset.
@@
Effect of $\theta$ (Lipschitz threshold) As shown in Fig 4(b) and Table VI, we analyze the effect of the training threshold $\theta$ (See Eq III). For smaller values of $\epsilon$, SPLITZ with higher Lipschitz constant achieves better performances on accuracy. Conversely, SPLITZ with a smaller Lipschitz constant can boost certified radius, which obtains a relative higher certified test accuracy when $\epsilon$ is larger. These observations suggest a trade-off between the Lipschitz constant and certified accuracy. Thus, identifying the optimal Lipschitz threshold is essential for achieving a balances between robustness and utility. This ablation study further validates that the essence of our SPLITZ classifier lies in maintaining a relatively optimized Lipschitz constant for the left half of the classifier.
Effect of $\gamma$ (size of radius around input $x$) According to above results, constraining the local Lipschitz constant achieves better performance. To further explore the benefit of local Lipschitz constrained training, it is necessary to explore the indicator matrix $I^{v}$ in Eq 18, which depends on the size of the ball around the input (i.e., the hyperparameter $\gamma$). In Table VII, we show how varying training $\gamma$ impacts the certified test accuracy for different values of $\epsilon$. We observe that smaller values of training $\gamma$ lead to higher certified accuracy for all values of $\epsilon$.

### v2

@@ IV Experiments and Evaluation Results
In this section, we evaluate the SPLITZ classifier on three datasets, MNIST[36], CIFAR-10 [37] and ImageNet [38]. We also present results on the Adult Income and Law school dataset in Appendix C, demonstrating that SPLITZ achieves better trade-offs between robustness and accuracy on both tabular and image datasets. We report the approximate certified test accuracy and certified radius of smoothed classifiers over full test datasets in MNIST and CIFAR-10 datasets and a subsample of 1,000 test data in ImageNet dataset. Same as previous works, we vary the noise level $\sigma\in\{0.25,0.5,1.0\}$ for the smoothed models on CIFAR-10 and ImageNet dataset, $\sigma\in\{0.25,0.5,0.75,1.0\}$ for MNIST dataset. We certified the same noise level $\sigma$ during the inference time. To ensure a fair comparison with previous works, we provide the highest reported results from each paper for the corresponding above levels of noise magnitudes. To improve certified robustness, we utilize the tighter local Lipschitz bound introduced in [19]. For three datasets, we use the same model as previous works [7, 14, 12, 22] (LeNet for MNIST, ResNet110 for CIFAR-10, ResNet50 for ImageNet). More experimental details are described in Appendix C.
@@
MNIST Certified Test Accuracy at $\epsilon~{}(\%)$
Method Extra data 1.50 1.75 2.00 2.25 2.50
RS [7] 67.3 46.2 32.5 19.7 10.9
MACER [16] 73.0 50.0 36.0 28.0 -
Consistency [22] 82.2 70.5 45.5 37.2 28.0
SmoothMix [12] 81.8 70.7 44.9 37.1 29.3
DRT [23] 83.3 69.6 48.3 40.3 34.8
RS+OrthoNN [21] 70.1 49.7 33.2 21.0 10.5
SPLITZ 80.2 71.3 62.3 51.7 40.5
TABLE II: Comparison of certified test accuracy (%) on MNIST under $\ell_{2}$ norm perturbation. Each entry lists the certified accuracy using numbers taken from respective papers (RS results follow from previous benchmark papers [22, 12]).
CIFAR-10 Certified accuracy at $\epsilon~{}(\%)$
@@
TABLE III: Comparison of the approximate certified test accuracy (%) on CIFAR-10 under $\ell_{2}$ norm perturbation. Extra data indicates whether their models incorporate other datasets in their models. Each entry lists the certified accuracy using numbers taken from respective papers.
SPLITZ Methodology For all datasets, we split the classifier after $1^{\text{st}}$ affine layer where the left half of the classifier contains one convolution layer followed by the clipped ReLU layer (See Appendix B). For the ImageNet dataset, the only difference is that we remove the BatchNorm layer after the $1^{st}$ affine layer and we replace the ReLU layer with the clipped ReLU layer in the first half of the network, which helps us obtain a tighter local Lipschitz bound of the first half of the classifier. The rest of the classifier is the same as original models (LeNet for MNIST, ResNet110 for CIFAR-10, ResNet50 for ImageNet).
Dataset Configuration For the MNIST and CIFAR-10 dataset(s), we draw $n_{0}=10^{2},~{}n_{1}=10^{5}$ noise samples to certify the smoothing model following [7, 14, 12]. For ImageNet dataset, we draw $n_{0}=10^{2},~{}n_{1}=10^{4}$ noise samples to certify the smoothing model following [7, 14, 12]. We set the Lipschitz threshold (see Sec III) as $\theta=0.5$. For local Lipschitz constrained training, we set tradeoff parameter $\lambda$ (see Sec III) evenly decrease from $0.8-0.4$, $0.7-0.5$ and $0.9-0.7$ respectively. We use one Nvidia P100 GPU to train the SPLITZ model with batch size $512,256,128$ respectively. We apply Adam Optimizer for three datasets. For the MNIST dataset, we conduct training for 150 epochs and utilize an early-stop strategy to search for the optimal classifier over an additional 150 epochs. We set the initial learning rate as 0.001. The learning rate is decayed (multiplied by 0.1) by 0.1 at every 50 epochs (50th, 100th…). For the CIFAR-10 dataset, we train 200 epochs using the ResNet110 and use the early-stop strategy to search for the optimal classifier over an additional 200 epochs. Furthermore, we set the initial learning rate as 0.001 and final learning rate as $10^{-6}$. The learning rate starts to evenly decay at each epoch from the half of the training epochs. For the ImageNet dataset, we train 200 epochs for the ResNet50 and set the initial learning rate as 0.01. The learning rate starts to decay at each 40 epochs. Our code is available at: https://anonymous.4open.science/r/SPLITZ-3C66. We report our more experimental details in Appendix C.
@@ IV-A Main Results
Results on MNIST As showed in Table II, we can observe that SPLITZ outperforms other state-of-art approaches in almost every value of $\epsilon$. Impressively, we find that the SPLITZ classifier has a significant improvement when the value of $\epsilon$ is large. For instance, when $\epsilon=2.50$, SPLITZ classifier achieves 40.5% compared to state-of-art top-1 test accuracy 34.8% certified test accuracy on the MNIST dataset. Moreover, when we increase $\epsilon$ from 1.50 to 2.50, RS drops from 67.3% to 10.9% decreasing 56.4% test accuracy. SPLITZ, however, maintains higher certified test accuracy from 80.2% to 40.5% maintaining relatively higher test accuracy.
@@
Results on CIFAR10 As shown in Table III and Fig 3, our method outperforms the state-of-art approaches for every value of $\epsilon$ on CIFAR-10 dataset. Interestingly, we find that the SPLITZ training has a significant improvement when the value of $\epsilon$ is large. For instance, when $\epsilon=1.0$, the model achieves 43.2% top-1 test accuracy on CIFAR-10 dataset compared to state-of-art top-1 test accuracy 39.8%. One hypothesis is that minimizing the Lipscitz bound of $f_{L}$ ($L_{f_{L}}\leq 1$) is able to boost the certified radius of the model. Intuitively, more samples are correctly classified while corresponding radius are larger than given $\epsilon$. In addition, we can observe the similar trend as MNIST dataset. SPLITZ maintains higher certified test accuracy when we increase $\epsilon$ from 0.25 to 1.00 compared to other state-of-art mechanisms.
ImageNet Certified accuracy at $\epsilon~{}(\%)$
Method Extra data 1 1.5 2.0 3.0
PixelDP [8] 0.0 0.0 0.0 0.0
RS [7] 37.0 29.0 19.0 12.0
SmoothAdv [15] 43.0 37.0 27.0 20.0
MACER [16] 43.0 31.0 25.0 14.0
Consistency [22] 44.0 34.0 24.0 17.0
SmoothMix [12] 43.0 38.0 26.0 20.0
Boosting [17] 44.6 38.4 28.6 21.2
DRT [23] 44.4 39.8 30.4 23.2
ACES [39] 42.2 35.6 25.6 19.8
DDS [14] 54.3 38.1 29.5 13.1
RS + OrthoNN [21] 38.0 28.4 16.2 10.0
SPLITZ (ours) 43.2 38.6 31.2 23.9
TABLE IV: Comparison of the approximate certified test accuracy (%) on ImageNet under $\ell_{2}$ norm perturbation. The columns and rows have the same meaning as in Table III.
@@
Results on ImageNet We show the comparison of different certified robustness techniques on ImageNet dataset in Table IV. We observe similar trends to MNIST and CIFAR10 datasets, where SPLITZ is effective on certified robustness with a wide range of image datasets. For instance, when $\epsilon=2.0$, SPLITZ achieve $31.2\%$ while the state-of-the-art have $30.4\%$. Moreover, SPLITZ consistently achieves better ACR (average certified radius) than other mechanisms in the ImageNet dataset, as shown in Table V.
@@
Results on Average Certified Radius (ACR) We investigate the performance of SPLITZ using average certified radius (ACR), where we measure the correct samples’ average certified radius over the test datasets (MNIST, CIFAR-10, ImageNet). As shown in Table V, we provide the comprehensive comparison results of average certified radius (ACR) compared to other certified robust techniques. Our SPLITZ consistently outperforms others. For instance, when $\sigma=0.50$, the ACR of SPLITZ is 2.059 respectively, where the state-of-the-art is 0.933 in MNIST dataset. In addition, we can observe the same trends in CIFAR-10 and ImageNet dataset, where SPLITZ consistently outperforms the state-of-the-art when $\sigma=0.50,1.00$.
Empirical Robust Test Accuracy We study the empirical robust test accuracy in Table VI under $\ell_{2}$ FGSM, PGD and Auto-attack using RS and SPLITZ model with variance $\sigma=1$. We can observe that the SPLITZ consistently outperforms the RS method. For example, when $\epsilon=2.0$ under auto-attack, SPLITZ achieve the empirical test accuracy $23.15\%$ while RS is $7.72\%$.
@@ IV-B Training and certifying time
Our SPLITZ model needs to vary the value of $\gamma$ (the size of the ball around input $x$) during the training epoch. Thus, we may need relatively more time to obtain the optimal model. To solve this, we apply the early stop mechanism to obtain a better optimized model during the training process. At the same time, we decay our learning rates during the training process. We use one Nvidia P100 GPU to train our SPLITZ model and report our training time (certifying time) for each one epoch (one image) in the Table VII.
$\sigma$ Methods MNIST CIFAR-10 ImageNet
0.50 $\text{RS}^{*}$ 1.553 0.525 0.733
$\text{SmoothAdv}^{*}$ 1.687 0.684 0.825
MACER 1.583 0.726 0.831
Consistency 1.697 0.726 0.822
SmoothMix 1.694 0.737 0.846
SPLITZ 2.059 0.924 0.968
1.00 $\text{RS}^{*}$ 1.620 0.542 0.875
SmoothAdv 1.779 0.660 1.040
MACER 1.520 0.792 1.008
Consistency 1.819 0.816 0.982
SmoothMix 1.823 0.773 1.047
SPLITZ 2.104 0.979 1.282
TABLE V: Comparison of average certified radius (ACR) across three different datasets (MNIST, CIFAR-10 and ImageNet). We can observe that for two datasets, SPLITZ consistently achieves better results compared to other state-of-art mechanisms. * is reported by [22, 12]
FGSM
Methods $\epsilon=0.5$ $\epsilon=1.0$ $\epsilon=2.0$
RS 46.86 38.79 25.54
SPLITZ 59.96 54.69 46.77
PGD
Methods $\epsilon=0.5$ $\epsilon=1.0$ $\epsilon=2.0$
RS 33.35 23.63 10.18
SPLITZ 50.53 39.05 27.03
Auto-attack
Methods $\epsilon=0.5$ $\epsilon=1.0$ $\epsilon=2.0$
RS 32.39 21.90 7.72
SPLITZ 49.35 33.21 23.15
TABLE VI: Comparison of empirical certified test accuracy under attacks with respect to CIFAR-10 dataset.
@@ IV-C Ablation Study
Training
Datasets Architecture Training per epoch (s) Certifying per image (s)
MNIST LeNet 15.1 1.0
CIFAR-10 ResNet110 59.5 50.7
ImageNet ResNet50 4810.5 83.1
TABLE VII: SPLITZ training time and certifying time across three datasets (MNIST using LeNet, CIFAR-10 using ResNet 110 and ImageNet using ResNet50).
Certified Test Accuracy at $\epsilon~{}(\%)$
Location of Splitting 0.50 0.75 1.00 1.25 1.50 1.75 2.00
$1^{\text{st}}$ affine layer 94.1 92.0 88.8 84.7 79.0 71.3 62.3
$2^{\text{nd}}$ affine layer 89.7 86.2 81.0 75.0 68.0 59.9 51.0
$3^{\text{rd}}$ affine layer 84.4 80.4 75.6 69.9 63.6 56.9 49.4
TABLE VIII: Comparison of certified test accuracy of SPLITZ with Gaussian noise $\sigma=0.75$ for varying the splitting layer on MNIST dataset with LeNet.
@@
Impact of splitting location As mentioned in Section III, our SPLITZ classifier can be optimized over different split ways, where we conduct the experiments and show our results in Table VIII. For example, when $\epsilon=2$, splitting after the $1^{st}$, or $2^{nd}$, or $3^{rd}$ layer result in certified accuracy of $62.3$%, $51.0$% and $49.4$% respectively. These results indicate that splitting the neural network early achieves better performance. Intuitively, splitting the neural network early helps the model minimize the local Lipschitz bound, which improves the certified robustness leading to a higher certified test accuracy given the same $\epsilon$. As the splitting becomes ”deeper”, estimating the local Lipshitz constant also becomes harder, which implies that a looser bound leads to smaller certified radius.
@@
Comparison of the orthogonal neural network (OrthoNN) and the SPLITZ In the baseline RS+OrthoNN model, during the training process, we enforce the first convolution layer of the classifier’s left half to act as an orthogonal convolution neural network. This is achieved by incorporating an orthogonal constraint regularization into the loss function. For the SPLITZ model, we impose a constraint on the local Lipschitz constant of the classifier’s left half. The comparison between RS+orthoNN and SPLITZ is illustrated in Table II with respect to the MNIST dataset, Table III with respect to the CIFAR-10 dataset and Table IV with respect to the ImageNet dataset. Our observations indicate that enforcing an orthogonal convolution layer on the classifier’s left half has a comparatively lesser impact than training with a local Lipschitz constraint.
@@
TABLE IX: Comparison of certified test accuracy of SPLITZ with Gaussian noise $\sigma=0.75$ for varying $Lip_{\theta}$ (the threshold of the local Lipschitz bound around input $x$) on MNIST dataset.
@@
TABLE X: Comparison of certified test accuracy of SPLITZ with Gaussian noise $\sigma=0.75$ for varying $\gamma$ (the size of the ball around input $x$) on MNIST dataset.
Effect of $\theta$ (Lipschitz threshold) As shown in Fig 4(b) and Table IX, we analyze the effect of the training threshold $\theta$ (See Eq III). For smaller values of $\epsilon$, SPLITZ with higher Lipschitz constant achieves better performances on accuracy. Conversely, SPLITZ with a smaller Lipschitz constant can boost certified radius, which obtains a relative higher certified test accuracy when $\epsilon$ is larger. These observations suggest a trade-off between the Lipschitz constant and certified accuracy. Thus, identifying the optimal Lipschitz threshold is essential for achieving a balances between robustness and utility. This ablation study further validates that the essence of our SPLITZ classifier lies in maintaining a relatively optimized Lipschitz constant for the left half of the classifier.
Effect of $\gamma$ (size of radius around input $x$) According to above results, constraining the local Lipschitz constant achieves better performance. To further explore the benefit of local Lipschitz constrained training, it is necessary to explore the indicator matrix $I^{v}$ in Eq 18, which depends on the size of the ball around the input (i.e., the hyperparameter $\gamma$). In Table X, we show how varying training $\gamma$ impacts the certified test accuracy for different values of $\epsilon$. We observe that smaller values of training $\gamma$ lead to higher certified accuracy for all values of $\epsilon$.
Figure 5: Left: Certified Radius of SPLITZ method vs the size of the ball around the input $\gamma$. Right: Local Lipschitz Constant vs the size of the ball around the input $\gamma$.
Effect of $\gamma$ and local Lipshitz constant In this section, we study the optimization of the choice of $\gamma$. We show the tradeoffs between the certified radius of SPLITZ and the value of $\gamma$ in Fig 5 (Left). We randomly select 10 datapoints from CIFAR-10 dataset and certify them using the SPLITZ model with variance of 0.25. We also show the tradeoffs between the local Lipschitz constant and the value of $\gamma$. We observe that both the local Lipschitz constant and the the certified radius stabilizes after a certain increase in $\gamma$. Therefore, in practice, a one-step optimization is sufficient to obtain the certified radius of SPLITZ. However, if the optimal certified radius is desired, a binary search can be performed, although it may be more time-consuming than one-step optimization.
IV-D Results on Tabular Dataset
Figure 6: Tradeoffs between the certified robustness and accuracy with respect to the RS [7], local Lipschitz constant constrained [19] and SPLITZ method on the Adult Income Dataset.
Figure 7: Tradeoffs between the certified robustness and accuracy with respect to the RS [7], local Lipschitz constant constrained [19] and SPLITZ method on the Law School Dataset.
Results on Adult Income Dataset Adult income dataset includes income related data with 14 features (i.e., age, work class, occupation, education etc.) of $N=45,222$ users ($N_{train}=32,561$, $N_{test}=12,661$) to predict whether the income of a person exceeds a threshold (e.g., \$50k) in a year.
We study the certified robustness of three robust classifiers (RS [7], Local Lipschitz constrained [19] and SPLITZ) on the Adult dataset shown in Fig 6. We observe that SPLITZ achieves better tradeoffs between the certified robustness and accuracy.
Results on Law School Dataset Law School dataset includes the admission related data with 7 features (LSAT score, gender, undergraduate GPA etc.) of $N=4,862$ applicants ($N_{train}=3,403$, $N_{test}=1,459$) to predict the likelihood of passing the bar.
Similar to the previous comparison, we evaluate SPLITZ against two other methods: Randomized Smoothing (RS) [7] and Local Lipschitz Constraints [19] as shown in Fig 7. Our findings demonstrate that SPLITZ consistently surpasses these techniques, achieving a significantly larger certified radius.